OpenStudy (anonymous):

Last week Talia walked, ran, and biked for a total of 7 hours. The number of miles she biked was two times the number of miles she walked. She walked 2 miles per hour, ran 6 miles per hour, and biked 10 miles per hour, traveling a total of 50 miles. How many hours did Talia run this week?

OpenStudy (abmon98):

W- Walking R - Running and B - Biking \[2W+6R+10B=50 \]W+R+B=7 \[B=2W \] \[3W+R=7 \] \[R=7-3W \] \[2W+6(7-3W)+10(2W)=50 \] \[2W+42-18W+20W=50 \] \[ 4W=8 \] \[W=2 \] \[B=2*2=4 \] \[R=7-6=1 \]